# Pragmatic Curiosity: A Hybrid Learning-Optimization Paradigm via Active Inference

### v1

@@ Appendix A Related Work
$D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]=\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],$
@@
$\begin{split}&D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]\\ =&D_{\text{KL}}[p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})]\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\int\frac{p(f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (6)
@@
$D_{\text{KL}}[P||Q]=\int_{\mathcal{X}}\log(\frac{dP}{dQ})dP,$
@@
$\begin{split}&D_{\text{KL}}[P(f_{\mathcal{X}})||\hat{P}(f_{\mathcal{X}})]\\ =&\int_{f_{\mathcal{X}}}\log(\frac{dP}{d\hat{P}}(f_{\mathcal{X}}))dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathcal{X}})\\ =&\int_{f_{\mathbf{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathbf{X}})\\ =&\mathbb{E}_{\hat{P}(f_{\mathbf{X}})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (7)
which has the exact same form as (6).
@@
According to Lemma C.1,
$\begin{split}I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D})=&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}[D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]]\\ =&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{L(\mathbf{Y})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\mathbb{E}_{L(\mathbf{Y})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]L(\mathbf{Y})d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int L(\mathbf{Y}|f_{\mathbf{X}})\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[D_{\text{KL}}[L(\mathbf{Y}|f_{\mathbf{X}})||L(\mathbf{Y})]]\\ =&I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}).\end{split}$ (8)
Appendix D Experimental Details
@@
D.1 Simulation Environment
All simulations for the perception failure evaluation in CARLA (Section D.3) were conducted on a Linux workstation with Ubuntu 22.04 LTS equipped with an Intel 13th Gen Core i7-13700KF CPU (16 cores, 24 threads, up to 5.4 GHz) and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM). The system ran CARLA simulations using CUDA 12.2 and NVIDIA driver version 535.230.02.
@@ D.1 Simulation Environment
D.2 Constrained System Identification
D.2.1 Plume Field Model and Parameters

### v2

@@ 6 Conclusion
$D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}}|\mathcal{D})]=\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],$
@@
$\begin{split}&D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}}|\mathcal{D})]\\ =&D_{\text{KL}}[p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})]\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\int\frac{p(f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (6)
@@
$D_{\text{KL}}[P\|Q]=\int_{\mathcal{X}}\log(\frac{dP}{dQ})dP,$
@@
$\begin{split}&D_{\text{KL}}[P(f_{\mathcal{X}})\|\hat{P}(f_{\mathcal{X}})]\\ =&\int_{f_{\mathcal{X}}}\log(\frac{dP}{d\hat{P}}(f_{\mathcal{X}}))dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathcal{X}})\\ =&\int_{f_{\mathbf{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathbf{X}})\\ =&\mathbb{E}_{\hat{P}(f_{\mathbf{X}})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (7)
which has the exact same form as equation 6.
@@
According to Lemma A.1,
$\begin{split}I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D})=&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}[D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}}|\mathcal{D})]]\\ =&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{L(\mathbf{Y})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\mathbb{E}_{L(\mathbf{Y})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]L(\mathbf{Y})d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int L(\mathbf{Y}|f_{\mathbf{X}})\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[D_{\text{KL}}[L(\mathbf{Y}|f_{\mathbf{X}})\|L(\mathbf{Y})]]\\ =&I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}).\end{split}$ (8)
Appendix B Re-interpretation of GP-UCB
The reinterpretations of most of the acquisition strategies in Table 1 are straightforward according to their definitions. However, placing a rather intuitive GP-UCB strategy within this framework seems implicit. To reveal their connection, we resort to Lemma 4.1. Then if we assume constant Gaussian noises $\mathcal{N}\sim(0,\sigma^{2})$ for the observations, we have
$\begin{split}I(f_{x};y|\mathcal{D}_{t})&=h(y\mid\mathcal{D}_{t})-H(y|f_{x},\mathcal{D}_{t})\\ &=\frac{1}{2}\log(1+\sigma^{-2}\sigma_{t}^{2}(x)),\end{split}$
If we choose $\beta=\frac{1}{2}\log(1+\sigma^{-2})$, then the epistemic term in PraC, i.e., $I(f_{x};y|\mathcal{D}_{t})$, provides an upper bound of the square of the exploration term $\beta^{1/2}\sigma_{t}(x)$ in GP-UCB.
This reveals the close relationship between GP-UCB and AIF, showing that the uncertainty bonus in GP-UCB can be viewed as a first-order surrogate for an information-theoretic epistemic term.
Appendix C Experimental Details
@@
C.1 Simulation Environment
All simulations for the perception failure evaluation in CARLA (Section C.3) were conducted on a Linux workstation with Ubuntu 22.04 LTS equipped with an Intel 13th Gen Core i7-13700KF CPU (16 cores, 24 threads, up to 5.4 GHz) and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM). The system ran CARLA simulations using CUDA 12.2 and NVIDIA driver version 535.230.02.
@@ C.1 Simulation Environment
C.2 Decision-Oriented Plume Monitoring
This appendix provides implementation details and additional analysis for the decision-oriented plume-monitoring experiments in Section 5.1. All three tasks share the same sequential decision structure. The agent maintains a posterior distribution over latent environmental hypotheses $\theta$, selects one sensor query $x$ at each iteration, observes a plume count $y$, updates the posterior, and evaluates the downstream decision induced by the updated belief.
C.2.1 Plume Field Model